OpenStudy (anonymous):

Find the equation of a line with the give properties: Through (6, 2) and perpendicular to the line with x-intercept 2 and y-intercept -3

OpenStudy (anonymous):

y = mx +c using 2 points the line, (x1, y1) and (x2, y2), where x2 > x1: \[m = \frac{y_2 - y_1}{x_2-x_1} \] c is value of y when x = 0. notice that the questions gives you 3 points, (6,2), (2,0) and (0,3). try solving using the method outlined above.

OpenStudy (anonymous):

@tridian I don't think so. the required line perpendicular to that line, it means dot product of 2 directions of two line is =0

OpenStudy (anonymous):

I think , we have to find out the equation of the line with x-intercept =2 and y-intercept =-3 to have it direction d_1

OpenStudy (anonymous):

then, use property of perpendicular vector to find out the direction d_2 of the required line from that direction + the given point, write down the equation of the line. slope line is just for the given line only.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

@Hoa oops, misread the question. thanks for spotting.